If hydrogen peroxide is your solute and saline is the solvent, to prepare 4 fl oz of 1/4 strength, how much hydrogen peroxide will you use?
Answer #1

Given data : Solute = hydrogen peroxide

Solvent = Saline

Required volume = 4 fl oz

Required strength = 1/4

We know that 1 L = 1000mL= 33.814 fl oz

Therefore 1 fl oz = 1000/33.814 = 29.573mL

So 4 fl oz = 29.573 ×4 = 118.29 mL

The solution contains 1/4 hydrogen peroxide and 3/4 saline.

Quantity of saline = 118.29 ×3/4 = 88.72mL

Quantity of hydrogen peroxide = 118.29 × 1/4 = 29.57mL